### v2

@@ Abstract
We provide a framework to study stability notions for two-sided dynamic matching markets in which matching is one-to-one and irreversible. The framework gives centerstage to the set of matchings an agent anticipates would ensue should they remain unmatched, which we refer to as the agent’s conjectures. A collection of conjectures, together with a pairwise stability and individual rationality requirement given the conjectures, defines a solution concept for the economy. We identify a sufficient condition–consistency–for a family of conjectures to lead to a nonempty solution (cf. Hafalir, 2008). As an application, we introduce two families of consistent conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability, and the extension to dynamic markets of the solution concept in Hafalir (2008), sophisticated dynamic stability.
@@ 1 Introduction
Formally, we study a two-sided, dynamic matching market, in which agents on both sides of the economy arrive over time. We assume arrivals are deterministic, that is, agents can perfectly foresee when each agent arrives in the economy.Footnote 2Footnote 2Footnote 2We assume deterministic arrivals because it simplifies notation, but it is not needed for our results. See 2.2. A matching specifies who matches with who in each period, subject to two constraints: agents cannot be matched before they arrive, and agents can have at most one matching partner on the opposite side. Agents are discounted utility maximizers, and their payoff from a matching solely depends on who they match with and when. Our goal is to study solution concepts for this economy.
@@
An agent’s conjectures describe the set of matchings the agent assumes may ensue should they remain unmatched in a given period. Importantly, a given conjectured matching describes both the current period’s matching outcome and the continuation matching from the next period onward. Because the set of agents available to match from the next period onward depends on both the newly arriving agents and those who remain unmatched from previous periods, the current period’s matching and the next period’s arrivals define the set of feasible continuation matchings. In other words, an agent’s outcome from remaining unmatched cannot be defined independently of other contemporaneous agents’ matching outcomes. Thus, having assumed away payoff externalities, an externality still exists in our model as other contemporaneous agents’ matching outcomes determine which continuation matchings are feasible. For a given conjectured matching, we can define an agent’s acceptable partners to be those preferred to the conjectured matching. Thus, to determine an agent’s acceptable partners, we need an assumption on how to select a continuation matching from the agent’s conjectures. Following Sasaki and Toda (1996), we say the agent finds a matching partner acceptable if the matching partner is preferred to some conjectured continuation matching.
@@
The paper’s main result, Theorem 4.1, provides a sufficient condition on the family of conjectures for matchings satisfying the solution concept to exist. We refer to the conjectures satisfying this property as consistent. Consistent conjectures guarantee that a candidate matching, which can be computed from the model primitives, is a solution for the economy (see Section 4 for a definition of the set of candidate matchings). Loosely speaking, consistency requires the following condition holds between the candidate matching–denote it by $m^{\star}$–and the conjectures. In each period and for each agent who is unmatched under $m^{\star}$ in that period, the candidate matching $m^{\star}$ should be one of the agent’s conjectures. In other words, when checking the continuation candidate matching against their own conjectures for what matchings would ensue given that they remain unmatched, the agents who are unmatched under $m^{\star}$ will find the candidate matching to be consistent with their conjectures. Consistency is related to the condition in Hafalir (2008, Proposition 2), which ensures the existence of stable matchings in static markets with payoff externalities; we explain the connection between the two in Section 4.
@@
Whereas consistency is defined relative to the candidate matching, an immediate corollary is that the following conjectures always lead to a nonempty solution concept. Suppose that each agent’s conjectures only rule out continuation matchings that do not conform with the solution concept in the continuation. For instance, in a two-period economy, these conjectures specify that in the last period only stable matchings among the remaining unmatched agents and the new arrivals are possible. These conjectures always satisfy consistency and Theorem 4.1 implies they lead to a non-empty solution concept. That is, if the agents conjecture that the economy conforms to the solution concept in the continuation and place no restrictions on the current period matching, the set of matchings that conform to the solution concept from today onward is nonempty.
@@
Theorem 4.1 suggests a way to define solution concepts for dynamic matching markets with nonempty solution correspondences by focusing attention on conjectures that satisfy consistency. We provide two illustrations of this approach in Section 5, where we present two such conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability and sophisticated dynamic stability (Hafalir, 2008). Continuation-value-respecting dynamic stability (henceforth, CVR-dynamic stability) is a refinement of dynamic stability (Doval, 2022). Like the conjectures in the previous paragraph, the conjectures in CVR-dynamic stability are defined recursively with the solution concept, with agents ruling out continuation matchings that are not $\mathrm{CVR}$-dynamically stable. In contrast to the conjectures in the previous paragraph, CVR-dynamic stability also places restrictions on the current period matching. In particular, each agent rules out that other contemporaneous agents leave the economy matched to matching partners that are worse than their worst conjectured matching partner. That is, each agent’s conjectures respect other agents’ continuation values as specified by their own conjectures. Sophisticated dynamic stability extends the solution concept in Hafalir (2008), sophisticated expectations, to dynamic matching markets. Sophisticated dynamic stability builds on the rational expectations solution, a classical solution concept in matching with externalities that does not have consistent conjectures (see 4.1). Building on the insight of Hafalir (2008), the conjectures of sophisticated dynamic stability are iteratively built from those corresponding to rational expectations to ensure their consistency.
@@
We are closest to the contribution of Hafalir (2008). In a static, two-sided, one-to-one matching market with payoff externalities, Hafalir (2008) provides a sufficient condition on agents’ conjectures under which stable matchings exist. As we discuss in detail in Section 4, Theorem 4.1 provides the analogue of Hafalir (2008, Proposition 2) for dynamic matching markets. Whereas conceptual differences between the papers exist, the logic behind the constructions is the same. Furthermore, we provide the natural extension of the solution concept in Hafalir (2008), sophisticated expectations, to the dynamic environments under consideration in Section 5.2. We also introduce a new solution concept for dynamic matching markets, CVR-dynamic stability, which has no analogue in his work.
@@
The literature on cooperative solution concepts for dynamic matching markets highlights the role of agents’ conjectures about what matching would ensue after a block in defining solution concepts (see, e.g., Kurino, 2009; Kadam and Kotowski, 2018; Kotowski, 2019; Altinok, 2019; Liu, 2023 for repeated matching markets and Doval, 2022; Nicolò, Salmaso, and Saulle, 2023, 2024 for irreversible matching, and Doval (Forthcoming) for a survey treatment of this literature). Because solution concepts in this literature are typically predicated on perfection requirements–i.e., the same solution concept applying at the appropriately defined continuation–agents’ conjectures are defined jointly with the solution concept. We follow a complementary approach in which we take agents’ conjectures as given and use them to define a solution concept. This allows us to identify a property of conjectures–consistency–that guarantees the corresponding solution concept is nonempty, and underlies existence results in the literature (cf. Doval, 2022). We also propose new solution concepts that refine those existing in the literature, illustrating how our result can be used as a recipe of sorts to define stability notions for dynamic matching markets.
@@
Recently, Nicolò, Salmaso, and Saulle (2024) study a third refinement of dynamic stability (Doval, 2022). Like the conjectures in dynamic stability, the conjectures in the solution concept in Nicolò et al. (2024) are defined recursively based on what the solution concept specifies in the continuation. In particular, the agents rule out all matchings that do not abide by the solution concept from tomorrow onward. Unlike dynamic stability, the solution concept in Nicolò et al. (2024) imposes more structure on what matchings the agents can conjecture will ensue in the current period when considering remaining unmatched: The authors define conjectures in a given period through an iterative procedure which, at each stage, rules out current period matchings that do not satisfy some minimal stability requirements given the conjectures at that stage.Footnote 3Footnote 3Footnote 3Whereas the main focus of the paper is on the conjectures that obtain as the limit of this procedure, by stopping the procedure at a given stage the authors can study stability under different levels of sophistication (Hafalir, 2008). It is possible to show the conjectures they construct satisfy consistency. Following Doval (2022), their proof is based on showing that (the analogue for their solution concept of) the candidate matchings we define in Section 4 satisfy the conditions of their solution concept. Furthermore, the authors show via an example that the rational expectations solution may fail to exist in a two-sided dynamic matching market.
@@ 2 Model
A length-$T$ economy, denoted $E^{T}=(A_{1},B_{1},\dots,A_{T},B_{T})$, consists of two sequences, $\{A_{1},\dots,A_{T}\}$ and $\{B_{1},\dots,B_{T}\}$, of disjoint subsets of $\mathscr{A}$ and $\mathscr{B}$, respectively.Footnote 4Footnote 4Footnote 4It is without loss to assume the sets $(A_{1},\dots,A_{T})$ are disjoint as we do not rule out indifferences. Thus, $a$ denotes a unique individual in the economy, but $a$ may be payoff-equivalent to other arriving agents on side $A$. The same applies for the sets $(B_{1},\dots,B_{T})$ being disjoint. For $t\in\{1,\dots,T\}$, $A_{t}$ are the side $A$-agents that arrive in period $t$, and $B_{t}$ are the side $B$-agents that arrive in period $t$. For any $t\leq T$, let $\overline{A}_{t}=\cup_{s=1}^{t}A_{s}$ denote the implied arrivals on side $A$ through period $t$; similarly, let $\overline{B}_{t}=\cup_{s=1}^{t}B_{s}$ denote the implied arrivals on side $B$ through period $t$.
@@
A period-$t$ matching is a mapping $m_{t}:\overline{A}_{t}\cup\overline{B}_{t}\mapsto\overline{A}_{t}\cup\overline{B}_{t}$ such that the following hold:
@@
For each $t\geq 1$, a matching $m$ induces a continuation economy of length $T-t$, $E_{t+1}^{T}(m^{t})=(\pazocal{A}_{t+1}(m^{t}),\pazocal{B}_{t+1}(m^{t}),\dots,A_{T},B_{T})$.
@@
We close the model by defining the agents’ preferences. Throughout, we assume agents discount the future. Formally, each $a\in\mathscr{A}$ defines a discount factor $\delta_{a}\in[0,1]$ and a Bernoulli utility, $u(a,\cdot):\mathscr{B}\cup\{a\}\mapsto\mathbb{R}$. Similarly, each $b\in\mathscr{B}$ defines a discount factor $\delta_{b}\in[0,1]$ and a Bernoulli utility, $v(\cdot,b):\mathscr{A}\cup\{b\}\mapsto\mathbb{R}$. We normalize to 00 the payoffs of remaining single for one period. That is, for all $a\in\mathscr{A}$ and all $b\in\mathscr{B}$, $u(a,a)=v(b,b)=0$.
@@
We note that several pieces of notation, like $\pazocal{M}_{T}$, $\pazocal{A}_{t}(\cdot)$, $\pazocal{B}_{t}(\cdot)$, denote objects whose definitions depend on the economy $E^{T}$ under consideration. To be precise, we should note them $\pazocal{M}_{T}(E^{T}),\pazocal{A}_{t}(\cdot;E^{T}),\pazocal{B}_{t}(\cdot,E^{T})$. In what follows, we continue to omit this dependence whenever it does not lead to confusion.
@@
We note two assumptions that simplify notation, but are otherwise not needed for the results. First, we assume agents can perfectly foresee who arrives in each period. Instead, our results extend to the case in which arrivals are stochastic so that an economy is defined as a distribution over sequences $E^{T}$ as in Doval (2021). Second, we assume time-discounted preferences, but all that matters for our results is that agents are dynamically consistent.
@@ 3 A solution concept given conjectures
Fix a period $t$, a matching through $t-1$, $m^{t-1}$, and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$. A conjecture for agent $k$ is a subset $\varphi_{t}(m^{t-1},k)$ of $\pazocal{M}_{T}(m^{t-1},k)$.
@@
In words, a matching $m$ is a $\Phi$-solution for $E^{T}$ if in each period the following holds. First, each agent $k$ who is available to match in period $t$ prefers $m$ to their worst conjectured matching (parts 1 and 2 in 6). Note that this implies that if $m_{t}(k)\neq k$, then $k$ prefers their matching partner to waiting to be matched, and if $m_{t}(k)=k$, then $k$ has no objections to the continuation matching, $(m_{s})_{s=t+1}^{T}$.Footnote 5Footnote 5Footnote 5In a static economy, only the first kind of block may exist. The second block only shows up in a dynamic economy (see Doval, 2022 for a discussion). For instance, if the continuation matching is not self-enforcing, an agent may object to it. As we discuss later on, consistency of the conjectures rules is very much related to the second class of blocks. In both cases, we say $k$ prefers to be available to match in period $t$. Second, no pair of agents who can match in period $t$ prefer to match with each other to matching according to $m$ (part 3 in 6). Note that $(a,b)$ being able to match in period $t$ does not imply $a$ and $b$ arrive in the same period, but that under matching $m$ they meet in period $t$ and are able to block $m$.
6 is in the spirit of the static notion of stability:Footnote 6Footnote 6Footnote 66 is also similar to that of dynamic stability in Doval (2022), except that the conjectures in dynamic stability are recursively defined using the solution concept. In other words, 6 nests that of dynamic stability by allowing for arbitrary conjectures. $\Phi$-solutions are defined by the absence of pairwise blocks and the requirement that each agent is matched to a partner that is preferred to remaining unmatched. In contrast to the static notion of stability, the value of remaining unmatched is determined by the conjectures. Indeed, when $T=1$, the correspondence $\Sigma_{1}^{\Phi}$ reduces to the set of stable matchings (4). To see this, consider a length-1 economy, $E^{1}=(A_{1},B_{1})$, a matching $m$ for $E^{1}$, and an agent $k\in A_{1}\cup B_{1}$. Because all matchings $\overline{m}\in\varphi_{1}(m^{0},k)$ are such that $\overline{m}_{1}(k)=k$, then parts 1 and 2 of 6 simply state that $k$ prefers $m$ to remaining single. That is, $m$ is individually rational. Furthermore, part 3 implies that $m$ has no pairwise blocks. Thus, when $T=1$, 6 reduces to 4 for any family of conjectures. We record this observation for future reference:
@@
Among other things, 1 implies that $\Phi$-solutions are always nonempty for economies of length $1$. We use this property repeatedly in our proofs.
Whereas 6 is fairly general, it is worth noting that the definition rules out solution concepts in which the conjectures depend on the matching under consideration beyond the induced economy in period $t$. For instance, when considering whether to block matching $m$ in period $t$, if agent $k$ presumes all other contemporaneous agents–except perhaps for $k$’s matching partner–remain matched according to $m_{t}$, then $k$’s conjectures depend on the outcome specified by $m$ in period $t$. (These conjectures are referred to as passive conjectures; see Chowdhury (2004) and Doval (2015).) Instead, 6 presumes we can define the conjectures without reference to any specific matching. The matching under consideration only matters inasmuch as it determines who can match in each period, and hence whose conjectures one should check in each period.
@@
    $(m_{t})|_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$.Footnote 7Footnote 7Footnote 7We abuse notation slightly. Whereas $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$ defines a matching only for the agents who can match from $t=2$ onward–i.e., it is an element of $\pazocal{M}_{T-1}(E_{2}^{T}(m_{1}))$$m$ also specifies the outcome for those who have matched in period $1$. Thus, part 4 should be read as “$(m_{t})_{t=2}^{T}$ coincides with an element of $\Sigma_{T-1}^{\Phi}(E^{T}_{2}(m_{1}))$ for the agents who are yet to be matched at the end of period 1, and coincides with $m_{1}$, otherwise.”
7 simply states that matching $m$ is a $\Phi$-solution for $E^{T}$ if the following two conditions hold. First, from period $2$ onwards, $m$ specifies a $\Phi$-solution for the induced economy of length $T-1$, $E_{2}^{T}(m_{1})$. Second, $m$ has no blocks in period $1$ in the sense of 6. Because the correspondence $\Sigma_{T-1}^{\Phi}$ only depends on the conjectures from period $2$ onward, this recursive definition allows us in what follows to focus on the properties of $\varphi_{1}(\emptyset,\cdot)$, taking as given that the set $\Sigma_{T-1}^{\Phi}$ is nonempty.

### v3

@@ Abstract
We provide a framework to study stability notions for two-sided dynamic matching markets in which matching is one-to-one and irreversible. The framework gives center stage to the set of matchings an agent anticipates would ensue should they remain unmatched, which we refer to as the agent’s conjectures. A collection of conjectures, together with a pairwise stability and individual rationality requirement given the conjectures, defines a solution concept for the economy. We identify a sufficient condition—consistency—for a family of conjectures to lead to a nonempty solution (cf. Hafalir, 2008). As an application, we introduce two families of consistent conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability, and the extension to dynamic markets of the solution concept in Hafalir (2008), sophisticated dynamic stability.
@@ 1 Introduction
Formally, we study a two-sided, dynamic matching market, in which agents on both sides of the economy arrive over time. We assume arrivals are deterministic; that is, agents can perfectly foresee when each agent arrives in the economy.Footnote 2Footnote 2Footnote 2We assume deterministic arrivals because doing so simplifies notation, but this assumption is not needed for our results. See 2.2. A matching specifies who matches with whom in each period, subject to two constraints: agents cannot be matched before they arrive, and agents can have at most one matching partner on the opposite side. Agents are discounted utility maximizers, and their payoff from a matching depends only on whom they match with and when. Our goal is to study solution concepts for this economy.
@@
An agent’s conjectures describe the set of matchings the agent assumes may ensue should they remain unmatched in a given period. Importantly, a given conjectured matching describes both the current-period’s matching outcome and the continuation matching from the next period onward. Because the set of agents available to match from the next period onward depends on both the newly arriving agents and those who remain unmatched from previous periods, the current-period’s matching and the next period’s arrivals define the set of feasible continuation matchings. In other words, an agent’s outcome from remaining unmatched cannot be defined independently of other contemporaneous agents’ matching outcomes. Thus, having assumed away payoff externalities, an externality still exists in our model because other contemporaneous agents’ matching outcomes determine which continuation matchings are feasible. For a given conjectured matching, we can define an agent’s acceptable partners as those preferred to the conjectured matching. Thus, to determine an agent’s acceptable partners, we need an assumption on how to select a continuation matching from the agent’s conjectures. Following Sasaki and Toda (1996), we say the agent finds a matching partner acceptable if the matching partner is preferred to some conjectured continuation matching.
@@
The paper’s main result, Theorem 4.1, provides a sufficient condition on the family of conjectures for matchings satisfying the solution concept to exist. We refer to the conjectures satisfying this property as consistent (9). Consistent conjectures guarantee that candidate matchings, which can be computed from the model primitives, are a solution for the economy (see Section 4 for a definition of the set of candidate matchings). Loosely speaking, consistency requires that each candidate matching—denote it by $m^{\star}$—and the conjectures satisfy the following condition. In each period and for each agent who is unmatched under $m^{\star}$ in that period, the candidate matching $m^{\star}$ should be one of the agent’s conjectures. In other words, when checking the continuation candidate matching against their own conjectures for what matchings would ensue given that they remain unmatched, the agents who are unmatched under $m^{\star}$ will find the candidate matching to be consistent with their conjectures. Consistency is related to the condition in Hafalir (2008, Proposition 2), which ensures the existence of stable matchings in static markets with payoff externalities; we explain the connection between the two in Section 4.
@@
Whereas consistency is defined relative to the candidate matching, an immediate corollary is that the following conjectures always lead to a nonempty solution concept. Suppose each agent’s conjectures only rule out continuation matchings that do not conform with the solution concept in the continuation. For instance, in a two-period economy, these conjectures specify that in the last period, only stable matchings among the remaining unmatched agents and the new arrivals are possible. These conjectures always satisfy consistency, and Theorem 4.1 implies they lead to a nonempty solution concept. That is, if the agents conjecture that the economy conforms to the solution concept in the continuation and place no restrictions on the current-period matching, the set of matchings that conform to the solution concept from today onward is nonempty.
@@
Theorem 4.1 suggests a way to define solution concepts for dynamic matching markets with nonempty solution correspondences by focusing attention on conjectures that satisfy consistency. We provide two illustrations of this approach in Section 5, where we present two such conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability and sophisticated dynamic stability (Hafalir, 2008). Continuation-value-respecting dynamic stability (henceforth, $\mathrm{CVR}$ dynamic stability) is a refinement of dynamic stability (Doval, 2022). Like the conjectures in the previous paragraph, the conjectures in $\mathrm{CVR}$ dynamic stability rule out continuation matchings that are not $\mathrm{CVR}$ dynamically stable. In contrast to the conjectures in the previous paragraph, $\mathrm{CVR}$ dynamic stability also places restrictions on the current-period matching. In particular, each agent rules out that other contemporaneous agents leave the economy matched to matching partners that are worse than their worst conjectured matching partner. That is, each agent’s conjectures respect other agents’ continuation values as specified by their own conjectures. Sophisticated dynamic stability extends the solution concept in Hafalir (2008), sophisticated expectations, to dynamic matching markets. Sophisticated dynamic stability builds on the rational expectations solution, a classical solution concept in matching with externalities that does not have consistent conjectures (see 4.1). Building on the insight of Hafalir (2008), the conjectures of sophisticated dynamic stability are iteratively built from those corresponding to rational expectations to ensure their consistency.
@@
We are closest to the contribution of Hafalir (2008), who studies stability in static, two-sided, one-to-one, matching markets with payoff externalities. Observing the negative results in Sasaki and Toda (1996) follow from conjectures being exogenously defined, Hafalir (2008) provides a sufficient condition on agents’ endogenous conjectures under which stable matchings exist. As we discuss in detail in Section 4, Theorem 4.1 provides the analogue of Hafalir (2008, Proposition 2) for dynamic matching markets. Whereas conceptual differences between the papers exist, the logic behind the constructions is the same. Furthermore, we provide the natural extension of the solution concept in Hafalir (2008), sophisticated expectations, to the dynamic environments under consideration in Section 5.2. We also introduce a new solution concept for dynamic matching markets, $\mathrm{CVR}$ dynamic stability, which has no analogue in his work.
In a static model of matching with contracts, Rostek and Yoder (2024) observe that implicit assumptions on agents’ beliefs, rather than payoff externalities or complementarities, are responsible for nonexistence results. The authors show stable matchings always exist when agents’ contract choices are derived from correct beliefs about the contracts others will choose. Similarly, in $\mathrm{CVR}$ dynamic stability, agents correctly anticipate others’ continuation values when deciding whether to block. However, they need not correctly anticipate others’ choices in the event of a block, and existence follows from the consistency of the conjectures.
@@
The literature on cooperative solution concepts for dynamic matching markets highlights the role of agents’ conjectures about what matching would ensue after a block in defining solution concepts (for repeated matching markets, see, e.g., Kurino, 2009; Kadam and Kotowski, 2018; Kotowski, 2019; Altinok, 2019; Liu, 2023; for irreversible matching, see, e.g., Doval, 2022; Nicolò, Salmaso, and Saulle, 2023, 2024, and for a survey treatment of this literature, see Doval (Forthcoming)). Because solution concepts in this literature are typically predicated on perfection requirements—that is, the same solution concept applying at the appropriately defined continuation—agents’ conjectures are defined jointly with the solution concept. We follow a complementary approach in which we take agents’ conjectures as given and use them to define a solution concept. This approach allows us to identify a property of conjectures—consistency—that guarantees the corresponding solution concept is nonempty, and underlies existence results in the literature (cf. Doval, 2022). We also propose new solution concepts that refine those in the existing literature, illustrating how our result can be used as a recipe to define stability notions for dynamic matching markets.
@@
Recently, Nicolò, Salmaso, and Saulle (2024) study a third refinement of dynamic stability (Doval, 2022). Like the conjectures in dynamic stability, the conjectures in the solution concept in Nicolò et al. (2024) are defined recursively based on what the solution concept specifies in the continuation. In particular, the agents rule out all matchings that do not abide by the solution concept from tomorrow onward. Unlike dynamic stability, the solution concept in Nicolò et al. (2024) imposes more structure on what matchings the agents can conjecture will ensue in the current period when considering remaining unmatched: the authors define conjectures in a given period through an iterative procedure that, at each stage, rules out current-period matchings that do not satisfy some minimal stability requirements given the conjectures at that stage.Footnote 3Footnote 3Footnote 3Whereas the main focus of the paper is on the conjectures that obtain as the limit of this procedure, by stopping the procedure at a given stage the authors can study stability under different levels of sophistication (Hafalir, 2008). The conjectures they construct satisfy consistency. Following Doval (2022), their existence proof is based on showing that (the analogue for their solution concept of) the candidate matchings we define in Section 4 satisfy the conditions of their solution concept. Furthermore, the authors show via an example that the rational expectations solution may fail to exist in a two-sided dynamic matching market.
Finally, our work is related to the literature on the farsighted stable set, which is used to model externalities in coalition-formation games (e.g., Harsanyi, 1974; Chwe, 1994; Ray and Vohra, 2015), and more recently, one-to-one matching markets (e.g., Mauleon et al., 2011; Kimya, 2022). As in this literature, agents in our model understand the terminal consequences of their moves. Whereas farsighted stability focuses on the credibility of coalitional blocks, the solution concepts we introduce in Section 5 focus on the credibility of the continuation matchings used to dissuade agents from blocking.
@@ 2 Model
A length-$T$ economy, denoted $E^{T}=(A_{1},B_{1},\dots,A_{T},B_{T})$, consists of two sequences, $\{A_{1},\dots,A_{T}\}$ and $\{B_{1},\dots,B_{T}\}$, of disjoint subsets of $\mathscr{A}$ and $\mathscr{B}$, respectively.Footnote 4Footnote 4Footnote 4Without loss, we can assume the sets $(A_{1},\dots,A_{T})$ are disjoint because we do not rule out indifferences. Thus, $a$ denotes a unique individual in the economy, but $a$ may be payoff equivalent to other arriving agents on side $A$. The same applies for the sets $(B_{1},\dots,B_{T})$ being disjoint. For $t\in\{1,\dots,T\}$, $A_{t}$ are the side-$A$ agents that arrive in period $t$, and $B_{t}$ are the side-$B$ agents that arrive in period $t$. For any $t\leq T$, let $\overline{A}_{t}=\cup_{s=1}^{t}A_{s}$ denote the implied arrivals on side $A$ through period $t$; similarly, let $\overline{B}_{t}=\cup_{s=1}^{t}B_{s}$ denote the implied arrivals on side $B$ through period $t$.
@@
A period-$t$ matching is a mapping $m_{t}:\overline{A}_{t}\cup\overline{B}_{t}\rightarrow\overline{A}_{t}\cup\overline{B}_{t}$ such that the following hold:
@@
For each $t\geq 1$, a matching $m$ induces a continuation economy of length $T-t$, $E_{t+1}^{T}(m^{t})=(\pazocal{A}_{t+1}(m^{t}),\pazocal{B}_{t+1}(m^{t}),\dots,A_{T},B_{T})$. Given a matching $\overline{m}\in\pazocal{M}_{T}(m^{t})$, we denote the restriction of $\overline{m}$ to the continuation economy $E_{t+1}^{T}(m^{t})$ by $(\overline{m}_{s})_{s=t+1}^{T}|_{E_{t+1}^{T}(m^{t})}$.
@@
We close the model by defining the agents’ preferences. Throughout, we assume agents discount the future. Formally, each $a\in\mathscr{A}$ defines a discount factor $\delta_{a}\in[0,1]$ and a Bernoulli utility, $u(a,\cdot):\mathscr{B}\cup\{a\}\rightarrow\mathbb{R}$. Similarly, each $b\in\mathscr{B}$ defines a discount factor $\delta_{b}\in[0,1]$ and a Bernoulli utility, $v(\cdot,b):\mathscr{A}\cup\{b\}\rightarrow\mathbb{R}$.Footnote 5Footnote 5Footnote 5We follow the convention that $10^{0}=$ 1. We normalize to 00 the payoffs of remaining single for one period. That is, for all $a\in\mathscr{A}$ and all $b\in\mathscr{B}$, $u(a,a)=v(b,b)=0$.
@@
We note several pieces of notation, such as $\pazocal{M}_{T}$, $\pazocal{M}_{T}(m^{t-1})$, $\pazocal{A}_{t}(\cdot)$, $\pazocal{B}_{t}(\cdot)$, denote objects whose definitions depend on the economy $E^{T}$ under consideration. To be precise, we should note them: $\pazocal{M}_{T}(E^{T}),\pazocal{M}_{T}(m^{t-1};E^{T}),\pazocal{A}_{t}(\cdot;E^{T}),\pazocal{B}_{t}(\cdot;E^{T})$. In what follows, we continue to omit this dependence whenever doing so does not lead to confusion.
@@
We note two assumptions that simplify notation, but are otherwise not needed for the results. First, we assume agents can perfectly foresee who arrives in each period. Our results extend to the case in which arrivals are stochastic so that an economy is defined as a distribution over sequences $E^{T}$ as in Doval (2021). Second, we assume time-discounted preferences, but all that matters for our results is that agents are dynamically consistent.
@@ 3 A solution concept given conjectures
Fix a period $t$, a matching through $t-1$, $m^{t-1}$, and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$. A conjecture for agent $k$ is a nonempty subset $\varphi_{t}(m^{t-1},k)$ of $\pazocal{M}_{T}(m^{t-1},k)$.
@@
In words, a matching $m$ is a $\Phi$-solution for $E^{T}$ if in each period the following holds. First, each agent $k$ who is available to match in period $t$ prefers $m$ to their worst conjectured matching (parts 1 and 2 in 6). Note this condition implies that if $m_{t}(k)\neq k$, $k$ prefers their matching partner to waiting to be matched, and if $m_{t}(k)=k$, $k$ has no objections to the continuation matching, $(m_{s})_{s=t+1}^{T}$.Footnote 6Footnote 6Footnote 6In a static economy, only the first kind of block may exist. The second block only shows up in a dynamic economy (see Doval, 2022, for a discussion). For instance, if the continuation matching is not self-enforcing, an agent may object to it. As we discuss later, consistency of the conjectures is related to the second class of blocks. In both cases, we say $k$ prefers to be available to match in period $t$. Second, no pair of agents who can match in period $t$ prefer to match with each other to matching according to $m$ (part 3 in 6). Note $(a,b)$ being able to match in period $t$ does not imply $a$ and $b$ arrive in the same period, but that under matching $m$, they meet in period $t$ and are able to block $m$.
6 is in the spirit of the static notion of stability: $\Phi$-solutions are defined by the absence of pairwise blocks and the requirement that each agent is matched to a partner that is preferred to remaining unmatched.Footnote 7Footnote 7Footnote 76 is also similar to that of dynamic stability in Doval (2022), except the conjectures in dynamic stability are recursively defined using the solution concept. In other words, 6 nests that of dynamic stability by allowing for arbitrary conjectures. In contrast to the static notion of stability, the value of remaining unmatched is determined by the conjectures. Indeed, when $T=1$, the correspondence $\Sigma_{1}^{\Phi}$ reduces to the set of stable matchings (4). To illustrate, consider a length-1 economy, $E^{1}=(A_{1},B_{1})$, a matching $m$ for $E^{1}$, and an agent $k\in A_{1}\cup B_{1}$. Because all matchings $\overline{m}\in\varphi_{1}(\emptyset,k)$ are such that $\overline{m}_{1}(k)=k$, parts 1 and 2 of 6 simply state that $k$ prefers $m$ to remaining single. That is, $m$ is individually rational. Furthermore, part 3 implies $m$ has no pairwise blocks. Thus, when $T=1$, 6 reduces to 4 for any family of conjectures. We record this observation for future reference:
@@
Among other things, 1 implies $\Phi$-solutions are always nonempty for economies of length $1$. We use this property repeatedly in our proofs.
Whereas 6 is fairly general, note the definition rules out solution concepts in which the conjectures depend on the matching under consideration beyond the induced economy in period $t$. For instance, when considering whether to block matching $m$ in period $t$, if agent $k$ presumes all other contemporaneous agents—except perhaps for $k$’s matching partner—remain matched according to $m_{t}$, $k$’s conjectures depend on the outcome specified by $m$ in period $t$. (These conjectures are referred to as passive conjectures; see Chowdhury (2004) and Doval (2015).) Instead, 6 presumes we can define the conjectures without reference to any specific matching. The matching under consideration only matters inasmuch as it determines who matches in each period, and hence whose conjectures one should check in each period.
@@
    $(m_{t})_{t=2}^{T}|_{E^{T}_{2}(m_{1})}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$.
7 simply states matching $m$ is a $\Phi$-solution for $E^{T}$ if the following two conditions hold. First, from period $2$ onwards, $m$ specifies a $\Phi$-solution for the induced economy of length $T-1$, $E_{2}^{T}(m_{1})$. Second, $m$ has no blocks in period $1$ in the sense of 6. Because the correspondence $\Sigma_{T-1}^{\Phi}$ depends only on the conjectures from period $2$ onward, this recursive definition allows us in what follows to focus on the properties of $\varphi_{1}(\emptyset,\cdot)$, taking as given that the set $\Sigma_{T-1}^{\Phi}$ is nonempty.